OpenStudy (anonymous):

Make "x" the subject of: 1. a + x = b 2. ax = b 3. 2x + a = d

OpenStudy (anonymous):

doesnt make much sense

OpenStudy (anonymous):

-.- thats why I'm here....it doesn't make sense to me.

OpenStudy (anonymous):

x = b - a x = b/a x = (d - a)/2

OpenStudy (anonymous):

Alright LA-MrC, So basically leaving "x" by itself, and transferring the others "variables" to the other side?

OpenStudy (anonymous):

"make x the subject of ..." is bad math teacherese for "solve for x"

OpenStudy (anonymous):

Yes. Now x is the "subject" of each equation (statement).

OpenStudy (anonymous):

Okay I appreciate it ! Thank you.

OpenStudy (anonymous):

what it really means is write \[x=\text{expression}\]